A boat on a river travels downstream (with the current) between two points 20 miles apart, in one hour. The return trip (against the current) takes 2& 1/2 hours. What is the boat's speed, and what is the speed of the current?
Answer by lwsshak3(11628)

A boat on a river travels downstream (with the current) between two points 20 miles apart, in one hour. The return trip (against the current) takes 2& 1/2 hours. What is the boat's speed, and what is the speed of the current?
***
let x=boat's speed (in still water)
let c=speed of current
(x+c)boat's speed (downstream)
(x+c)boat's speed (against current)
distance/speed=travel time
..
20/(x+c)=1
20/x-c=2.5
..
x+c=20
2.5x-2.5c=20
..
2.5x+2.5c=50
2.5x-2.5c=20
add
5x=70
x=14
c=20-x=6
boat's speed (in still water)=14 mph
let c=speed of current=6 mph